Fermi National Laboratory


Accelerator Update

Friday July 16
The day shift began with CDF and D0 in access, and stacking in the Antiproton Source (Pbar).

Tevatron (Tev) shot set began at 9:29 AM.

Store #3657 was colliding at 11:52 AM with an initial luminosity of 103.43 E30. This is new record initial luminosity.

At 12:50 PM beam was established to MTest.

PBar transfers to the Recycler Ring (RR) started at 3:12 PM and were completed by 4:16 PM. We then returned to stacking.

Saturday July 17
The Midnight shift began with store #3657 colliding, stacking, and with beam to MiniBoone and Mtest.

At 1:05 AM MTest turned off for the night and MCenter (MIPP) turned on.

MTest turns on at 10:32 AM and MCenter turns off at 11:04 AM.

Store # 3657 was terminated at 8:01 PM.

Shot setup began at 8:45 PM, with store #3659 colliding at 11:00 PM with an initial luminosity of 30.45 E30. Since luminosity was so low we began investigating what went wrong with this shot. It appears the low luminosity is related to a problem with the Damper.

Pbar returns to stacking at 11:08 PM.

Sunday July 18
The Midnight shift began with store #3659 colliding, stacking, and with beam to MiniBoone.

Booster RF station #5 (BRF5) went down on the midnight shift and was disconnected from the Anode Supply. Repair may require an access into Booster.

Main Injector (MI) RF #12 went down also on the midnight shift and was disconnected from the anode supply. Repair will require an access into the main injector.

Store #3659 was terminated at 3:00 PM

Tev shot setup began at 3:48 PM, it appears that the damper is working correctly.

Store #3661 was colliding at 6:08 PM with an initial luminosity of 75.41 E30.

Booster RF station #2 (BRF2) tripped off at 7:40 PM, expert was contacted, and is coming in to look at the station.

At 10:00 PM we decided to run without BRF2 and repair it in the morning, we are currently down 2 Booster RF stations.

At 11:21 PM we began shot setup to the Recycler Ring, this was completed by 12:26 AM.

Monday July 19
The Midnight shift began with store #3361 colliding, stacking, and with beam to MiniBoone. Pbar stash in the recycler is 41 E10.

Today's Plans
Repair BRF5 and BRF2, and continue to stack Pbars.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/19/2004   email Fermilab